# Structure Your Data: Towards Semantic Graph Counterfactuals

### v1

@@ 4.3 Extendability of graph-based counterfactuals
We replicate Dervakos et al. (2023)’s experiment on explaining the classification of web-crawled creative-commons images into ’driver’ and ’pedestrian’ classes. Here, images were manually classified by the authors; thus, we explain a non-neural classifier. By employing the SoTA scene graph generator (SGG) of Cong et al. (2023) we extract global edits from generated graphs for the transition from ’pedestrian’ to ’driver’ (Fig. 5(left)). Their relevance is verified by our common sense: people wear helmets when driving -addition of (helmet, on, head) and (man, on, bike)- and cover the bike seat with their body -deletion of (seat, on, bike)-. To validate our method’s consistency across other annotation techniques, we replace the SGG with a pipeline of captioning (BLIP (Li et al., 2022)) and graph parsing (Unified VSE (Wu et al., 2019)). We confirm that resulting edits (Fig. 5 (right)) semantically resemble the ones in Fig. 5 (left). Overall, similarly to Fig. 1, more accurate local edits are achieved through the consideration of the multiplicity of objects and relations. Generic triple edits result from errors in the automatic annotation pipeline, emphasizing the importance of meticulous explanation dataset curation.
@@
Despite focusing on images, we briefly demonstrate our method’s model-agnostic nature by applying it for audio classification, following SC. We provide CEs using the Smarty4covid dataset (Zarkogianni et al., 2023) for the IEEE COVID-19 sensor informatics competition winner [^2] IEEE COVID-19 sensor informatics competition, which predicts COVID-19 from cough audio. Our findings align with SC, revealing the high frequency of concept edits among respiratory symptoms and uncovering the same gender bias (§G). No new findings are produced for this primarily concept-based dataset with trivial interconnections, once again placing the focus on the nature and density of the annotations. However, we confirm that our method is at least as good as SC in these cases nonetheless.
Figure 5: Graph edits (triples inserted/ deleted) to implement the ’pedestrian’ $\rightarrow$ ’driver’ transition.
@@ 4.4 Efficiency of graph-based counterfactuals
[…]
CUB 46220 32691 0.033 0.060
VG-DENSE 13982 12059 0.033 0.063
VG-RANDOM 18787 16271 0.029 0.099
Time Performance for Counterfactual Retrieval
We experimentally confirm that our method allows for efficient CE retrieval: In Tab. 8, we report execution times for CE computation on the complete sets of graphs using GED (Fankhauser et al., 2011) versus our GNN-powered approach. We further report retrieval and inference time of our method. Even by adding times for all GCN-N/2 operations, we significantly relieve the computational burden of calculating the ground truth GED for all graph pairs, especially for larger graphs.
@@
In Fig. 6, we examine how retrieval precision varies using different numbers of training pairs $p$. The consistency of behavior exhibited over 70K pairs concludes our claim that N/2 training pairs are adequate for appropriate graph embedding using GCN.
Figure 6: P@k of GCN variant for different training pairs $p$.
@@ Appendix B Graph statistics
VG-DENSE
VG-RANDOM
CUB
D/P-SGG
D/P-CAPTION
SMARTY
Mean density
0.199
0.061
0.037
0.128
0.247
0.234
edges
9.042
8.768
27.519
9.367
1.756
4.398
nodes
7.252
14.566
28.519
9.733
3.197
5.398
isolated nodes
0.47
3.37
0
0.323
0.901
0
Max density
0.467
0.667
0.111
1.0
0.5
0.333
edges
36
27
53
18
4
15
nodes
15
20
54
18
5
16
isolated nodes
3
12
0
3
4
0
Min density
0.144
0.013
0.019
0.046
0.05
0.062
edges
5
5
8
1
1
2
nodes
6
4
9
2
2
3
isolated nodes
0
0
0
0
0
0
@@
AG
GQA
Mean density
0.189
0.241
edges
13.227
8.144
nodes
8.84
6.66
isolated nodes
1.173
1.366
Max density
0.45
1.0
edges
51
20
nodes
17
12
isolated nodes
2
15
Min density
0.1
0.125
edges
4
5
nodes
5
4
isolated nodes
0
0
In Tab. 9 we present some additional statistics regarding the graphs of the datasets used in our work (max and min details). VG-DENSE and VG-RANDOM contain 500 graphs each, CUB contains 422 graphs, D/P-SGG and D/P-CAPTION denote the web-crawled datasets with 259 graphs each and SMARTY denotes the COVID-19 classification dataset with 548 graphs. Table 10 contains additional statistics about datasets utilized only in the appendix. These are GQA (Hudson & Manning, 2019) with 500 graphs mentioned in §D.3 and Action Genome (AG) (Ji et al., 2020) with 300 graphs mention in §F. The size and density of input data should be considered when viewing results in the experimental section.
@@ E.1 Counterfactual graph geometry on CUB
Our framework is capable of retrieving counterfactual graphs that not only respect node and edge semantics, but also graph geometry. This observation corresponds to more accurate retrieval capabilities that focus on semantic information regarding bird species without being significantly distracted from irrelevant characteristics such as the background. This can be an encouraging characteristic of our counterfactuals towards more robust explanations, even though this aspect is not analyzed in the current paper. First, we present a qualitative example of this claim. In Figure 17, we search for the most similar image to 16(a) using the method of CVE and ours.
@@
Apparently, both counterfactual images are visually similar, as appearing in Fig. 16(b) and 16(c). However, the representation power of scene graphs becomes evident in this case. In Fig. 18 we present the scene graphs corresponding one-to-one to the images of Fig. 17. The most similar graphs of 17(a) correspond to the graph of 17(b) according to CVE and 17(c) according to our approach. It is evident that our approach can successfully retrieve graphs that better respect the geometry of the source image scene graph. Another observation is that our approach manages to retrieve an image without the concepts ’leg’ or ’tail’ which is more accurate compared to the source. Therefore, structural similarity leads to better semantic consistency.
@@ E.3 Actionability of edits
(Vandenhende et al., 2022) often propose single edits on the query image (left images of Figs. 20(a), 20(b)) and deem them sufficient for the transition from query to target class. However, as explained in Sec. 4.1 of the main paper, this approach disregards the rest of the edits needed to be made between $I_{(A)}$ and $I_{(B)}$ and leads to instances that are in fact out-of-distribution. In the main paper, we gave an example that corresponds to Fig. 20(a). In addition to the combination (’has_head_pattern::eyering’, ’has_breast_color::grey’) that was reported in text, we provide several other attribute combinations that do not exist in any other bird of the target class in Tab. 16.
@@ Appendix G Applicability on other modalities
Triple Edits Norm. Counts
@@
Concept Edits Norm. Counts

### v2

@@ 4.3 Extendability of graph-based counterfactuals
We replicate Dervakos et al. (2023)’s experiment on explaining the classification of web-crawled creative-commons images into ’driver’ and ’pedestrian’ classes. Here, images were manually classified by the authors; thus, we explain a non-neural classifier. By employing the SotA scene graph generator (SGG) of Cong et al. (2023) we extract global edits from generated graphs for the transition from ’pedestrian’ to ’driver’ (Fig. 5(left)). Their relevance is verified by our common sense: people wear helmets when driving -addition of (helmet, on, head) and (man, on, bike)- and cover the bike seat with their body -deletion of (seat, on, bike)-. To validate our method’s consistency across other annotation techniques, we replace the SGG with a pipeline of captioning (BLIP (Li et al., 2022)) and graph parsing (Unified VSE (Wu et al., 2019)). We confirm that resulting edits (Fig. 5 (right)) semantically resemble the ones in Fig. 5 (left). Overall, similarly to Fig. 1, more accurate local edits are achieved through the consideration of the multiplicity of objects and relations. Generic triple edits result from errors in the automatic annotation pipeline, emphasizing the importance of meticulous explanation dataset curation. We provide further details regarding the ’pedestrian’ vs ’driver’ classification experiment in App. F. Additionally, we experiment with more unannotated datasets of scene images, such as Action Genome (Ji et al., 2020), which is presented in App. F.
@@
Despite focusing on images, we briefly demonstrate our method’s model-agnostic nature by applying it for audio classification, following SC. We provide CEs using the Smarty4covid dataset (Zarkogianni et al., 2023) for the IEEE COVID-19 sensor informatics competition winner [^2] IEEE COVID-19 sensor informatics competition, which predicts COVID-19 from cough audio. Our findings align with SC, revealing the high frequency of concept edits among respiratory symptoms and uncovering the same gender bias. No new findings are produced for this primarily concept-based dataset with trivial interconnections, once again placing the focus on the nature and density of the annotations. However, we confirm that our method is at least as good as SC in these cases nonetheless. More results regarding the audio classification experiment are discussed in App. G.
Figure 5: Graph edits (triples inserted/ deleted) to implement the ’pedestrian’ $\rightarrow$ ’driver’ transition. The yellow color distinguishes edge from node labels within a triple.
@@ 4.4 Efficiency of graph-based counterfactuals
Time Performance for Counterfactual Retrieval
From a theoretical standpoint, our method’s efficiency is expected. The heaviest part of GNN computations occurs during model training, where each backward call is correlated with the square of the number of nodes in addition to the number of edges. Inference, on the other hand, is nearly instantaneous and is linearly correlated with the sum of the number of nodes and edges in a graph (Blakely et al., 2019).
Additionally, we experimentally confirm that our method allows for efficient CE retrieval: In Tab. 8, we report execution times for CE computation on the complete sets of graphs using GED (Fankhauser et al., 2011) versus our GNN-powered approach. We further report retrieval and inference time of our method. Even by adding times for all GCN-N/2 operations, we significantly relieve the computational burden of calculating the ground truth GED for all graph pairs, especially for larger graphs.
@@
GED $\downarrow$ GCN-N/2 (train)$\downarrow$ GCN-N/2 (retr.$\downarrow$ GCN-N/2 (infer.$\downarrow$
CUB 46220 32691 0.03 0.06
VG-DENSE 13982 12059 0.03 0.06
VG-RANDOM 18787 16271 0.03 0.10
@@
In Fig. 6, we examine how retrieval precision varies using different numbers of training pairs $p$ on CUB (Fig. 6(b)) and the two VG variants (Fig. 6(a)). P@k does not exhibit significant increase in any case after the $\sim N/2$ pairs mark (70k for VG and 50k for CUB). On the contrary, it could remain identical (Fig. 6(a) left) or even decrease (Fig. 6(b)). The same precision pattern per $p$ is experimentally validated on the GQA dataset (Appendix D.3). The consistency of behavior exhibited over $\sim N/2$ pairs concludes our claim that $N/2$ training pairs are adequate for appropriate graph embedding using GCN.
(a) VG variants
(b) CUB
Figure 6: P@k of GCN variant for different training pairs $p$ on the two main datasets explored.
@@ 5 Conclusion
Acknowledgments
This work has been developed as part of the HiDALGO2 project, which has received funding from the European High Performance Computing Joint Undertaking (JU) and Poland, Germany, Spain, Hungary, France and Greece under grant agreement No 101093457. Maria Lymperaiou was supported by the Hellenic Foundation for Research and Innovation (HFRI) under the 3rd Call for HFRI PhD Fellowships (Fellowship Number 5537). We thank all reviewers for their insightful comments and feedback, and our annotators for participating in the conducted human surveys.
@@ Appendix B Graph statistics
VG-DENSE VG-RANDOM CUB D/P-SGG D/P-CAPTION SMARTY
Mean density 0.20 0.06 0.04 0.13 0.25 0.23
edges 9.04 8.77 27.52 9.37 1.76 4.40
nodes 7.25 14.57 28.52 9.73 3.20 5.40
isolated nodes 0.47 3.37 0 0.32 0.90 0
Max density 0.47 0.67 0.11 1.0 0.5 0.33
edges 36 27 53 18 4 15
nodes 15 20 54 18 5 16
isolated nodes 3 12 0 3 4 0
Min density 0.14 0.01 0.02 0.05 0.05 0.06
edges 5 5 8 1 1 2
nodes 6 4 9 2 2 3
isolated nodes 0 0 0 0 0 0
@@
AG GQA
Mean density 0.19 0.24
edges 13.23 8.14
nodes 8.84 6.66
isolated nodes 1.17 1.37
Max density 0.45 1.0
edges 51 20
nodes 17 12
isolated nodes 2 15
Min density 0.1 0.13
edges 4 5
nodes 5 4
isolated nodes 0 0
In Tab. 9 we present some additional statistics regarding the graphs of the datasets used in our work (max and min details). VG-DENSE and VG-RANDOM contain 500 graphs each, CUB contains 422 graphs, D/P-SGG and D/P-CAPTION denote the web-crawled datasets with 259 graphs each and SMARTY denotes the COVID-19 classification dataset with 548 graphs. Table 10 contains additional statistics about datasets utilized only in the appendix. These are GQA (Hudson & Manning, 2019) with 500 graphs mentioned in App. refsec:quant-additional and Action Genome (AG) (Ji et al., 2020) with 300 graphs mention in App. refsec:unannotated. The size and density of input data should be considered when viewing results in the experimental section.
@@ E.1 Counterfactual graph geometry on CUB
Our framework is capable of retrieving counterfactual graphs that not only respect node and edge semantics, but also graph geometry. This observation corresponds to more accurate retrieval capabilities that focus on semantic information regarding bird species without being significantly distracted from irrelevant characteristics such as the background. This can be an encouraging characteristic of our counterfactuals towards more robust explanations, even though this aspect is not analyzed in the current paper. First, we present a qualitative example of this claim. In Figure 17, we search for the most similar image to 17(a) using the method of CVE and ours.
@@
Apparently, both counterfactual images are visually similar, as appearing in Fig. 17(b) and 17(c). However, the representation power of scene graphs becomes evident in this case. In Fig. 18 we present the scene graphs corresponding one-to-one to the images of Fig. 17. The most similar graphs of 18(a) correspond to the graph of 18(b) according to CVE and 18(c) according to our approach. It is evident that our approach can successfully retrieve graphs that better respect the geometry of the source image scene graph. Another observation is that our approach manages to retrieve an image without the concepts ’leg’ or ’tail’ which is more accurate compared to the source. Therefore, structural similarity leads to better semantic consistency.
@@ E.3 Actionability of edits
(Vandenhende et al., 2022) often propose single edits on the query image (left images of Figs. 21(a), 21(b)) and deem them sufficient for the transition from query to target class. However, as explained in Sec. 4.1 of the main paper, this approach disregards the rest of the edits needed to be made between $I_{(A)}$ and $I_{(B)}$ and leads to instances that are in fact out-of-distribution. In the main paper, we gave an example that corresponds to Fig. 21(a). In addition to the combination (’has_head_pattern::eyering’, ’has_breast_color::grey’) that was reported in text, we provide several other attribute combinations that do not exist in any other bird of the target class in Tab. 16.
@@ Appendix G Applicability on other modalities
Concept Edits Normalized Counts
@@
Triple Edits Normalized Counts